# Adapting Large Language Models for Character-based Augmentative and Alternative Communication

### v1

@@ Abstract
Users of Augmentative and Alternative Communication (AAC) may write letter-by-letter via an interface that uses a character language model. However, most state-of-the-art large pretrained language models predict subword tokens of variable length. We investigate how to practically use such models to make accurate and efficient character predictions. We fine-tune models using a large dataset of sentences we curated in which each sentence is rated according to how useful it might be for spoken or written AAC communication. We find that using an algorithm to produce character predictions from a subword large language model provides more accurate predictions than adding a classification layer or using a byte-level model. We also find that our domain adaptation curriculum is effective at improving model performance on simple, conversational text.
@@ 2.1 Language Modeling for AAC Devices
Language models have been frequently studied in AAC applications in an attempt to improve communication rates. One way language models have been applied to AAC devices is through word completions and predictions. Word completions attempt to predict a user’s intended current word before they finish typing it. Word predictions attempt to predict subsequent words. Kristensson and Müllners (2021) show that word predictions can increase the text entry rate of users, but they can also decrease entry rate, depending on how and when predictions are displayed and the user’s typing strategy. Yusufali et al. (2023, 2024) used LLMs to generate word predictions for AAC devices and showed entry rates up to 30.4 WPM.
@@ 2.2 Applications for Character Predictions
While these approaches can help accelerate input with some AAC devices, other systems such as brain-computer interfaces (BCIs) rely on noisier signals. P300 spellers Farwell and Donchin (1988) present visual stimuli to users and measure their response using EEG electrodes. One example is the RSVP Keyboard Orhan et al. (2012), which shows a rapid series of characters to try and determine a user’s intended character. Using RSVP, it can take several series worth of user responses to gain enough evidence to type a character.
RSVP Keyboard uses an n-gram language model to generate a character-level probability distribution, which it combines with the EEG evidence to reduce the number of series required to select a character. We hypothesize a LLM will provide more accurate probabilities and help reduce the number of series needed to write a character.
Nomon Broderick and MacKay (2009); Bonaker et al. (2022b, a) allows users to select words or characters by activating a switch when the rotating clock corresponding to their target reaches noon. Similar to RSVP Keyboard, it can take several clicks to make a single selection, since the system needs to differentiate between similar clocks. Since Nomon provides both word and character selection options, it needs a language model that can evaluate the likelihood of each possibility to aid the user in selecting their target.
@@ 3 Character prediction algorithm
We then ask the model to estimate the likelihood of the next token and evaluate any token that matches our context. For efficiency, we only track a certain number of hypotheses at a time, known as the beam width. We continue expanding each hypothesis until it has regenerated the removed context tokens plus at least one character. We then store the likelihood for each final prediction in a list based on the character that directly follows the context. We continue this process until there are no more hypotheses to extend, or, for efficiency, the number of completed hypotheses exceeds a max completed parameter. Then we sum the likelihoods stored for each character in our symbol set and normalize to sum to one, arriving at a distribution over our symbol set. See out supplementary materials for an implementation of our search algorithm.
@@ 4.1 Private AAC-like test data
Unfortunately, actual messages from AAC users are difficult to source due to practical, ethical, and privacy reasons. There are a number of datasets that have been used to approximate AAC communications such as voice banking messages Costello (2014), Switchboard Godfrey et al. (1992), DailyDialog Li et al. (2017), COMM2 Vertanen (2013), and Turk dialogues Vertanen (2017). However, all of these datasets are available online and may have been used in training LLMs.
@@
We split each genre in half, creating a written dev set (8.6 K words), a written test set (9.0 K words), a spoken dev set (11.3 K words), and a spoken test set (10.8 K words). We will use the dev sets for most of our experiments, reserving the test sets for a final evaluation (Section 7). Our final evaluation will also use text written by an AAC user.
@@ 4.2 Perplexity experiments
As a strong n-gram baseline optimized for AAC-like text, we compare against a large character 12-gram language model [^2] https://osf.io/ajm7t/wiki/lm_char_dec19/. This model was trained using Witten-Bell smoothing on 21 B characters from Common Crawl, Reddit, movie subtitles, and Twitter. Training sentences were selected via cross-entropy difference selection Moore and Lewis (2010) using in-domain language models trained on DailyDialog Li et al. (2017), short email messages, and crowdsourced AAC-like messages Vertanen and Kristensson (2011).
@@
AAC 12-gram 70 2.547 0.1
distilgpt2 82 2.705 37.5
gpt2 124 2.540 42.2
gpt2-medium 355 2.448 52.6
gpt2-xl 1558 2.377 74.2
opt-125m 125 2.387 46.6
opt-350m 331 2.315 61.1
opt-1.3b 1316 2.220 60.9
opt-2.7b 2652 2.208 70.3
opt-6.7b 6658 2.198 70.5
opt-13b 12853 2.226 91.7
Meta-Llama-3.1-8B 8030 2.312 93.4
@@
4.3 Impact of left context case
Our previous perplexity results were on lowercase versions of our test sentences. Our search algorithms sums over all matching following tokens ignoring case. However, we conditioned character predictions on text context that was all lowercase. This simulates an AAC interface where the user does not bother to case their communications, or when casing is added later in some way. To measure the impact of lowercase-only context, we compared this approach against two other methods:
  • Human – We used the case entered by the worker and corrected by the authors. This serves as an upper bound on the gains possible with perfect automatic casing or explicit case input by the user.
  • Simple – We uppercased the first letter of the sentence and the first person singular subject pronouns: I, I’m, I’ll, I’ve, and I’d.
Using opt-350m, the perplexity on our dev set using all lowercase left context was 2.315. If instead, we used the human mixed case left context, perplexity was slightly lower at 2.299. The simple approach performed similarly at 2.303.
We think this simple approach was effective as 94% of uppercase words in the spoken dev set were either at the start of the sentence or involved the word “I”. Of mid-sentence words in uppercase, 41% did not involve the word “I” (e.g. days of the week), so there could be room to improve the simple heuristic. We will use the simple method for the remainder of our experiments.
@@ 5.1 In-domain datasets
To approximate written AAC, we used sentences from online forum posts made by people using a mobile device Vertanen and Kristensson (2021). We denote this dataset as Forum. Our rationale for choosing this dataset was that it encompasses a rich set of topics while also having a bias towards a more concise writing style.
For the spoken in-domain class, we used a combination of the DailyDialog corpus Li et al. (2017) and the BOLT SMS/Chat corpus Song et al. (2014). We denote these datasets as Daily and BOLT respectively. Our rationale for choosing these datasets is that while the text was actually written and not spoken, it focuses on person-to-person dialogues. Compared to other datasets based on spoken dialogues such as Switchboard Godfrey et al. (1992), the text is relatively clean and without spoken speech artifacts such as filled pauses, false starts, or repetitions. Daily had the smallest amount of data, with 46 K sentences (356 K words). BOLT was the next largest at 271 K sentences (1.9 M words) followed by Forum at 888 K sentences (10.1 M words).
@@ 5.2 Large datasets
We filtered the full training corpora to extract only sentences that started with a capital letter, had both uppercase and lowercase characters, and ended with sentence end punctuation. We dropped sentences where more than 20% of the words were out-of-vocabulary with respect to a large list of 860 K words obtained from human-edited dictionaries. Throughout this work, we trained and tested on isolated sentences. We removed repeated identical sentences in Subtitle. We reserved 5% of the data as a dev set and 5% as a test set. After filtering, our training sets had 87.5 B words (C4) and 528 M words (Subtitle).
@@ 5.3 Sentence classification
Daily was the smallest with 46 K sentences. We used NLPAug Ma (2019) to add 232 K sentences, changing one or two words (selected at random) in the original sentences. Words were changed using a distilled version [^4] https://huggingface.co/distilroberta-base of RoBERTa Liu et al. (2019). This brought Daily to the same size (279 K sentences) as BOLT. We took the first 558 K sentences from Forum and News to create an equal amount of data for each of our three classes (in-domain written, in-domain spoken, and out-of-domain).
For purposes of classification, we lowercased all sentences and removed all characters aside from A–Z, apostrophe, and space. We optimized hyperparameters as described in Section B.1. We used early stopping during training. Our final model had a classification accuracy of 85.7% on unseen sentences. We ran our model on all sentences in C4 and Subtitle. We have made our classification model [^5] https://huggingface.co/figmtu/deberta-v3-aac-classifier and datasets of our scored sentences [^6] https://huggingface.co/datasets/figmtu/aac_c4_deberta_classified, [^7] https://huggingface.co/datasets/figmtu/aac_subtitle_deberta_classified available. These derivative datasets should only be used for research purposes.
@@ 5.4 Domain adaptation procedure
We used the scored C4 and Subtitle sentences to adapt the LLM. We selected sentences where either the probability of the written or spoken class was greater than or equal to a threshold. We trained 12-gram models from scratch and found we obtained the majority of perplexity gains with a threshold of 0.90 for C4 and 0.75 for Subtitle. Using these thresholds, our adaptation datasets consisted of 4.0 B words (C4) and 130 M words (Subtitle).
We explored a curriculum adaptation process consisting of a sequence of separate fine-tuning runs. Each run was on one of our five datasets: C4 (threshold 0.90), Subtitle (threshold 0.75), Daily, BOLT, or Forum. Before each fine-tuning, we optimized hyperparameters using the model from the previous step. See Section B.2 for details on our optimization and resulting hyperparameter values.
Given the sizes of our five datasets, we opted to first always fine-tune on C4 followed by Subtitle. We tested all possible orders of further fine-tuning on Daily, BOLT, and Forum. We found fine-tuning on BOLT followed by Daily performed the best. Fine-tuning on Forum did not help and thus we excluded it from the curriculum.
Written (ppl) Spoken (ppl) Avg (ppl)
opt-350m base 2.415 2.192 2.303
+ C4 0.90 2.342 2.091 2.217
      + Subtitle 0.75 2.242 2.091 2.167
        + BOLT 2.168 2.056 2.112
          + Daily 2.169 2.033 2.101
+ C4 random 2.468 2.168 2.318
+ All four 2.308 2.078 2.193
+ BOLT+Daily 2.265 2.104 2.184
Table 2: Performance of opt-350m on our two dev sets. The top section shows performance after each step of our domain adaptation curriculum. The bottom section shows three single-step adaptation approaches.
As shown in Table 2 (top), each step of the fine-tuning curriculum lowered the perplexity on one or both of our written and spoken dev sets. Overall, we reduced the perplexity of the original opt-350m model from 2.303 to 2.101 (8.8% relative). The final model is available on Hugging Face [^8] https://huggingface.co/figmtu/opt-350m-aac.
We compared our curriculum process against various other approaches (Table 2 bottom). To show that selecting sentences using our classification model helped, we fine-tuned on an equivalent random amount of C4. This model had a perplexity of 2.318, similar to the unadapted model. Rather than a multi-step fine-tuning curriculum, we could instead combine datasets and fine-tune just once. Single step fine-tuning on all four datasets improved perplexity to 2.193. Single step fine-tuning on just BOLT and Daily did slightly better at 2.184. This was a 5.2% relative reduction in perplexity compared to the unadapted model.
We also explored various other ways to try and improve our curriculum adaptation process. For details of things that did not work, see Section B.3.
@@ 6 Other prediction approaches
Our search algorithm (Section 3) takes an average of 66 ms to predict the next character on a A100 GPU. While this may be fast enough for many purposes, interfaces such as Dasher Ward and MacKay (2002) or less capable computing devices may require more efficient prediction. We explored several other methods for obtaining character predictions from a LLM.
@@ 6.1 Byte model
The multilingual encoder-decoder ByT5 model Xue et al. (2022) eschewed the use of subword tokenization and instead used a token set based on individual characters. ByGPT5 Belouadi and Eger (2023) is a decoder-only version of ByT5. Using ByGPT5, it is possible to make character predictions directly from the model’s output. We used the same fine-tuning curriculum process on ByGPT5 (289 M parameters) as we used on the subword opt-350m model.
The average perplexity on our dev sets of the base ByGPT5 model was 2.858. Adapating on our C4 classified data reduced perplexity substantially to 2.231. Likely this was due to the multilingual nature of the base model. Improvements diminished for subsequent fine-tuning steps. Our final model had a perplexity of 2.194, somewhat higher than our best opt-350m result of 2.101. This model is available on Hugging Face [^9] https://huggingface.co/figmtu/bygpt5-aac. Fine-tuning ByGPT5 was four times slower than opt-350m. However, predictions were over five times faster at 12 ms per character compared to 66 ms for opt-350m. See Section B.4 for per-step results and Section B.5 for hyperparameter values.
6.2 Classification
We also explored making character predictions by adding a classification layer to a transformer model. Instead of asking the model to generate the ensuing text given some context, we asked it to produce a distribution over classification labels – each label corresponded to a single character in our symbol set. All candidate classification models were built on top of the best domain-adapted opt-350m model from Section 5.4. This allowed us to retain the domain knowledge of the model and focus additional training on the classification task.
To create our training examples, we selected a random location within each sentence in the dataset. We truncated the sentence at that point, and labeled the example with the label corresponding to the first truncated character (i.e. the character that came next after the truncated text). If the next character was not in our symbol set (A–Z, a–z, space, and apostrophe), we chose another random location in the sentence. To train the classification layer of the model, we utilized the same dataset order and general procedure as the subword and byte models. For the smaller BOLT and Daily sets, we added a hyperparameter for the number of locations in each sentence that would be selected, creating multiple training examples from each sentence.
The classifier model did not perform as well as we had hoped, achieving a best perplexity of 2.338 on our dev sets by combining BOLT and Daily into a single training phase. This model is available on Hugging Face [^10] https://huggingface.co/figmtu/opt-350m-aac-character-classifier. By comparison, training with BOLT and then Daily in separate phases yielded a slightly higher perplexity of 2.352. See Section B.6 for the full table of results and Section B.7 for the hyperparameter values used for each step in the curriculum.
During our initial classifier experiments, we found that we could obtain slightly lower perplexities (on average around 0.01 lower) by adding a custom prefix to each training example and inference, in place of the model’s typical sentence start token. The results presented in this section used a prefix of “ AAC|”. We theorize that this method was effective because it helped set the classification training examples apart from the examples it had previously seen and signaled to the model there was a change in the task being performed.
Predictions using the classification model were about four times faster than the search algorithm, taking 16 ms on average. However, it did add additional steps to the fine-tuning process since the classification model was built on top of the domain adapted model. Additionally, the perplexity of the model increased compared to the search algorithm.
7 Final evaluation
Language model Domain adaptation Character predictions Written (ppl) Spoken (ppl) AAC (ppl) Time (ms)
@@ 7 Final evaluation
[…]
At this point, we have extensively used our private dev sets to guide our curriculum adaptation process and our choice of prediction algorithm. We conducted a final evaluation using our private test sets as well as a test set consisting of the actual text of an AAC user. The AAC test set consists of 3.6 K sentences (65 K words) written by a user of Dasher Ward and MacKay (2002). The text was from various public and private presentations given by the user.
As shown in Table 3, our curriculum adapted version of opt-350m using the search algorithm performed the best on the written (2.151) and spoken (1.999) private test sets. Simply fine-tuning opt-350m on BOLT+Daily yielded the lowest perplexity on the test set from the AAC user (2.242), though the base opt350-m (2.244) and curriculum adapted opt-350m (2.292) had similar perplexities.
Averaging the perplexity across all models, there was a noticeable increase in perplexity between our written (2.327) and spoken (2.202) test sets and the presentations given by the AAC user (2.525). While we hoped our data selection approach using diverse text from mobile forum messages would allow our adapted models to perform well on both simple everyday communications and more complex planned written communications, at least for this one user, this did not work. As a measure of text complexity, we calculated the number of characters per sentence. The AAC user’s sentences were on average 94.4 characters per sentence compared to the much shorter sentences in the written (31.4) and spoken (39.9) test sets.
Consistent with previous results, the byte LLM provided the fastest performance of all the LLMs and was 5.2 times faster than our search algorithm and 1.3 times faster than using a classification head.

### v2

@@ Abstract
Users of Augmentative and Alternative Communication (AAC) may write letter-by-letter via an interface that uses a character language model. However, most state-of-the-art large pretrained language models predict subword tokens of variable length. We investigate how to practically use such models to make accurate and efficient character predictions. We fine-tune models using a large dataset of sentences we curated in which each sentence is rated according to how useful it might be for spoken or written AAC communication. We find that using an algorithm to produce character predictions from a subword large language model provides more accurate predictions than adding a classification layer or using a byte-level model. We also find that our domain adaptation procedure is effective at improving model performance on simple, conversational text.
@@ 1 Introduction
We have made our algorithm’s source code, classified data sets, our classification model, and our best language models available. [^1] https://osf.io/ajm7t/
This paper is structured as follows: first, we motivate why character predictions are important for AAC text entry (Section 2) and describe our algorithm for making such predictions using a subword tokenized LLM (Section 3). Next, we find the best starting model, testing performance on a private collection of AAC-like text (Section 4). We then explore adapting the LLM using small in-domain sets and using text classified from much larger out-of-domain sets (Section 5). Next, we compare our search algorithm against other possible approaches (Section 6). Finally, we show our results on five test sets (Section 7) before wrapping up with discussion (Section 8) and conclusions (Section 9).
@@ 2.1 Language modeling for AAC devices
Language models have been frequently studied in AAC applications in an attempt to improve communication rates through, for example, word completions and predictions. Word completions attempt to predict a user’s intended current word before they finish typing it. Word predictions attempt to predict subsequent words. Kristensson and Müllners (2021) show that word predictions can increase or decrease a user’s entry rate depending on how and when predictions are displayed and the user’s typing strategy. Yusufali et al. (2023, 2024) used LLMs to generate word predictions for AAC devices and showed entry rates up to 30.4 WPM.
@@ 2.2 Applications for character predictions
While these approaches can help accelerate input with some AAC devices, other systems such as brain-computer interfaces (BCIs) rely on noisier signals. P300 spellers Farwell and Donchin (1988) present visual stimuli to users and measure their response using EEG electrodes. One example is the RSVP Keyboard Orhan et al. (2012), which shows a rapid series of characters to try and determine a user’s intended character. Using RSVP, it can take several series worth of user responses to gain enough evidence to type a character. RSVP uses a character n-gram language model in combination with the EEG evidence to reduce the number of series needed to select a character. We hypothesize a LLM will provide more accurate probabilities and reduce the number of series needed.
Nomon Broderick and MacKay (2009); Bonaker et al. (2022b, a) allows users to select words or characters by activating a switch when the rotating clock corresponding to their target reaches noon. It can take several clicks to make a single selection since the system needs to differentiate between similar clocks. Since Nomon provides both word and character selection options, it needs a language model that can evaluate the likelihood of each possibility to help accelerate writing.
@@ 3 Character prediction algorithm
We then ask the model to estimate the likelihood of the next token and evaluate any token that matches our context. For efficiency, we only track a certain number of hypotheses at a time, known as the beam width. We continue expanding each hypothesis until it has regenerated the removed context tokens plus at least one character. We then store the likelihood for each final prediction in a list based on the character that directly follows the context. We continue this process until there are no more hypotheses to extend, or, for efficiency, the number of completed hypotheses exceeds a max completed parameter. Then we sum the likelihoods stored for each character in our symbol set and normalize to sum to one, arriving at a distribution over our symbol set. See Appendix C for our pseudocode search algorithm or our supplementary materials for a Python implementation.
@@ 4.1 Private AAC-like test data
Unfortunately, actual messages from AAC users are difficult to source due to practical, ethical, and privacy reasons. There are a number of datasets that have been used to approximate AAC communications such as voice banking messages Costello (2014a), Switchboard Godfrey et al. (1992), DailyDialog Li et al. (2017), COMM2 Vertanen (2013), and Turk dialogues Vertanen (2017). However, all of these datasets are available online and may have been used in training LLMs.
@@
We split each genre in half, creating a written dev set (8.6 K words), a written test set (9.0 K words), a spoken dev set (11.3 K words), and a spoken test set (10.8 K words). We will use the dev sets for most of our experiments, reserving the test sets for a final evaluation (Section 7). Our final evaluation will also use text written by an AAC user, messages written by people with ALS for voice banking Costello (2014b), and responses to common conversational situations Vertanen (2013).
@@ 4.2 Perplexity experiments
We compare against a baseline of a 12-gram language model optimized for AAC-like text, denoted AAC n-gram. [^3] https://osf.io/c6mnz This model was trained using Witten-Bell smoothing on 21 B characters from Common Crawl, Reddit, movie subtitles, and Twitter. Training sentences were selected via cross-entropy difference selection Moore and Lewis (2010) using in-domain models trained on DailyDialog Li et al. (2017), short emails, and AAC-like messages Vertanen and Kristensson (2011).
@@
AAC n-gram 70 2.55 0.1
distilgpt2 82 2.71 37.5
gpt2 124 2.54 42.2
gpt2-medium 355 2.45 52.6
gpt2-xl 1558 2.38 74.2
opt-125m 125 2.39 46.6
opt-350m 331 2.32 61.1
opt-1.3b 1316 2.22 60.9
opt-2.7b 2652 2.21 70.3
opt-6.7b 6658 2.20 70.5
opt-13b 12853 2.23 91.7
Meta-Llama-3.1-8B 8030 2.31 93.4
@@ 5 Optimizing language models for AAC
Our results in Table 1 were on lowercase versions of our test sentences. Our search algorithm sums over all matching following tokens ignoring case. However, we conditioned predictions on all lowercase context. This simulates an AAC interface where the user does not bother to denote case, or when casing is added later in some way. We found we could slightly improve perplexity on our dev sets using opt-350m from 2.32 to 2.30 by upper casing the first letter of the sentence and the words: I, I’m, I’ll, I’ve, and I’d. We found this simple approach performed similar to using the human-supplied case. We will use this automatic casing method for the remainder of our experiments.
@@ 5.1 In-domain datasets
To approximate written AAC, we used sentences from forum posts made by people using a mobile device Vertanen and Kristensson (2021). We denote this dataset as Forum. Our rationale for using this dataset was that it contains a rich set of topics while also having a bias towards concise writing.
To approximate spoken AAC, we used the DailyDialog corpus Li et al. (2017) and the BOLT SMS/Chat corpus Song et al. (2014). We denote these datasets as Daily and BOLT respectively. Our rationale for choosing these datasets was that while the text was actually written and not spoken, it focuses on person-to-person dialogues. Compared to other datasets based on spoken dialogues such as Switchboard Godfrey et al. (1992), the text is relatively clean and does not contain spoken speech artifacts. Daily had 46 K sentences (356 K words), BOLT had 271 K sentences (1.9 M words), and Forum had 888 K sentences (10.1 M words).
@@ 5.2 Large datasets
We filtered C4 and Subtitle to sentences that started with a capital letter, had both uppercase and lowercase characters, and ended with sentence end punctuation. We dropped sentences where more than 20% of the words were out-of-vocabulary with respect to a list of 860 K words obtained from human-edited dictionaries. Throughout this work, we trained and tested on isolated sentences. We removed repeated identical sentences in Subtitle. We reserved 5% of the data as a dev set and 5% as a test set. After filtering, our C4 training set had 87.5 B words and Subtitle had 528 M words.
@@ 5.3 Sentence classification
Daily was the smallest with 46 K sentences. We used NLPAug Ma (2019) to add 232 K sentences, changing one or two words (selected at random) in the original sentences. Words were changed using a distilled version of RoBERTa Liu et al. (2019). This brought Daily to the same size (279 K sentences) as BOLT. We took the first 558 K sentences from Forum and News to create an equal amount of data for each of our three classes (in-domain written, in-domain spoken, and out-of-domain). During classification, we lowercased sentences and removed characters aside from A–Z, apostrophe, and space. We optimized hyperparameters as described in Section B.1. We used early stopping during training. Our final model had a classification accuracy of 85.7% on unseen sentences.
@@ 5.4 Domain adaptation procedure
We used the scored C4 and Subtitle sentences to adapt the LLM. We selected sentences with a written or spoken probability greater than or equal to a threshold. To efficiently determine which threshold to use, we trained 12-gram models from scratch across a broad range of thresholds and found we obtained the majority of perplexity gains with a threshold of 0.90 for C4 and 0.75 for Subtitle. Using these thresholds, our adaptation datasets consisted of 4.0 B words of C4 and 130 M words of Subtitle. With these thresholds as a starting point, we then trained opt-350m models with a narrower range of thresholds centered on 0.90 for C4 and 0.75 for Subtitle. The results, detailed in Section B.3, confirmed the thresholds we had found.
We first tested fine-tuning all the model’s parameters on a combined adaptation set consisting of C4 (threshold 0.90), Subtitle (threshold 0.75), BOLT, and Daily. As shown in Table 2, this reduced the perplexity on our dev sets from 2.30 to 2.19. Next, we explored a curriculum adaptation process consisting of a sequence of separate fine-tuning runs. Each run was on one of our five datasets: C4, Subtitle, Daily, BOLT, or Forum. Before each fine-tuning, we optimized hyperparameters using the model from the previous step.
Given the sizes of our five datasets, we opted to first always fine-tune on C4 followed by Subtitle. We tested all possible orders of further fine-tuning on Daily, BOLT, and Forum. Fine-tuning on Forum did not help in any order, and thus we excluded it from the curriculum. We found fine-tuning on C4, Subtitle, BOLT, and then Daily was the best order, improving perplexity to 2.10. Combining BOLT and Daily in a single final step did slightly better at 2.09. However, simply fine-tuning using the combined BOLT and Daily data did slightly better still at 2.07. While the classified C4 and Subtitle datasets did not prove useful for improving the subword LLM, they will be useful in improving other models (see Section 6).
Adaptation data Fine-tuning Perplexity
type steps Written Spoken   Avg
2.41 2.19 2.30
C4+Sub+BOLT+Daily full 1 2.31 2.08 2.19
[…]
BOLT+Daily full 1 2.11 2.04 2.07
[…]
BOLT+Daily LoRA 1 2.13 2.05 2.09
[…]
Additionally, we fine-tuned opt-350m using Low-Rank Adaptation (LoRA) Hu et al. (2021), which drastically reduces the number of trainable parameters for downstream tasks. As shown in Table 2 (bottom), LoRA yields slightly higher perplexities compared to full parameter fine-tuning. As such, we continued to use full fine-tuning for the remainder of our experiments.
For additional results and hyperparameter values, see Section B.2. We also explored other ways to try and improve adaptation. For details of things that did not work, see Section B.3.
@@ 6 Other prediction approaches
Our search algorithm (Section 3) takes on average 66 ms to predict the next character on a A100 GPU. While this may be fast enough for many purposes, interfaces such as Dasher Ward and MacKay (2002) or less capable devices may require more efficient prediction. We explored several other methods for obtaining faster predictions.
6.1 N-gram model
Inference with an n-gram model was very fast at 0.1 ms (Table 1). To serve as a baseline, and to provide an open training data n-gram model, we created a character 12-gram n-gram mixture model from C4 (threshold 0.90) and Subtitle (threshold 0.75) using SRILM Stolcke (2002). We created separate C4 and Subtitle language models using Witten-Bell smoothing and no count cutoffs. We built a mixture model using linear interpolation with weights optimized to minimize perplexity on dev sentences from Forum, BOLT, and Daily. The mixture weights were C4 0.84 and Subtitle 0.16.
The unpruned mixture model (denoted C4+Sub n-gram full) was large with 5 B parameters and a compressed size of 30 GB. We used entropy pruning Stolcke (1998) to to reduce the model to 71 M parameters and a compressed size of 488 MB (denoted C4+Sub n-gram). This was similar in size to the AAC n-gram we previously compared with (Table 1). On our dev sets, our full and pruned models had perplexities of 2.53 and 2.60 respectively. Our pruned model’s perplexity was somewhat higher than the 2.55 of the existing AAC n-gram model.
@@ 6.2 Byte model
The multilingual encoder-decoder ByT5 model Xue et al. (2022) eschewed the use of subword tokenization and instead used a token set based on individual characters. ByGPT5 Belouadi and Eger (2023) is a decoder-only version of ByT5 that can make character predictions directly.
The average perplexity on our dev sets of the base ByGPT5 model was 2.86. We compared our two best adaptation approaches found on opt-350m (Section 5.4). First, we used a three step curriculum approach, first adapting on C4, followed by Subtitle, and finally on BOLT combined with Daily. This yielded a model with a perplexity of 2.19.
Second, we fine-tuned in a single-step on BOLT combined with Daily. This results in a higher perplexity of 2.31. We suspect due to the multilingual nature of the base model, first fine-tuning on C4 and Subtitle helped move the model towards English. For results and hyperparameter values at each phase of the fine-tuning, see Section B.4.
Language Domain Character Written Spoken AAC user ALS bank COMM2 Time
model adaptation prediction (ppl) (ppl) (ppl) (ppl) (ppl) (ms)
AAC n-gram - KenLM 2.56 $\pm$ .03 2.51 $\pm$ .03 2.89 $\pm$ .01 2.33 $\pm$ .03 2.41 $\pm$ .02 0.1
C4+Sub n-gram - KenLM 2.62 $\pm$ .03 2.54 $\pm$ .02 2.86 $\pm$ .01 2.43 $\pm$ .03 2.45 $\pm$ .02 0.1
C4+Sub n-gram full - KenLM 2.57 $\pm$ .03 2.47 $\pm$ .02 2.82 $\pm$ .01 2.38 $\pm$ .03 2.40 $\pm$ .02 0.1
opt-350m - search alg. 2.40 $\pm$ .03 2.14 $\pm$ .02 2.24 $\pm$ .01 2.38 $\pm$ .03 2.20 $\pm$ .02 61.0
[…]
opt-350m BOLT+Daily search alg. 2.10 $\pm$ .02 2.00 $\pm$ .02 2.29 $\pm$ .01 2.13 $\pm$ .03 2.01 $\pm$ .02 61.6
@@
6.3 Classification
We also explored making character predictions by adding a classification layer to the LLM. Instead of asking the model to generate the ensuing text given some context, we asked it to produce a distribution over classification labels – each label corresponded to a single character in our symbol set. We built the model on top of the best domain-adapted opt-350m model from Section 5.4. This allowed us to retain the model’s domain knowledge and focus additional training on the classification task.
To create training examples, we selected a random location within each sentence. We truncated the sentence at that point and used the character that would come next as the label. If the next character was not in our symbol set (A–Z, a–z, space, and apostrophe), we chose another random location. We utilized the same three step curriculum approach as the byte model; we first trained on C4, then Subtitle, and finally on BOLT combined with Daily. For the smaller BOLT and Daily sets, we added a hyperparameter for the number training examples generated per sentence. During training, we updated the classification layer as well as all other model parameters.
The classifier model did not perform as well as we had hoped, achieving a perplexity of 2.47. By comparison, training with BOLT and then Daily in separate steps yielded a slightly higher perplexity of 2.54. Training on just BOLT and Daily performed worse at 2.60. See Section B.5 for additional results and hyperparameter values.
7 Final evaluation
At this point, we have extensively used our private dev sets to guide our process. We conducted a final evaluation with our unseen private test sets and three additional test sets. Firstly, we created a test set based on the text of an AAC user. This set consisted of 3.6 K sentences (65 K words) written by a user of Dasher Ward and MacKay (2002). The text was from public and private presentations given by the user. We also added two public test sets: voice banked messages of people with ALS Costello (2014a), and the COMM2 set of responses to conversational situations Vertanen (2013).
As shown in Table 3 (top), our C4+Sub n-gram model performed slightly worse with an average test set perplexity of 2.58 compared to the AAC n-gram model which averaged 2.54. Our unpruned full model did slightly better at 2.53. All n-gram models were outperformed by the LLMs, both with and without domain adaptation.
As shown in Table 3 (middle), using the search algorithm with opt-350m adapted in a single step on the combination of BOLT and Daily performed the best on four of the five test sets. It tied with the three- and four-step curriculum method on the spoken test set (2.00) and with the four-step curriculum on the COMM2 set (2.01). The base opt-350m yielded the lowest perplexity on the test set from the AAC user (2.24), though the curriculum adapted opt-350m models were not far behind (2.29 – 2.30).
Averaging the perplexity across all models, there was a noticeable increase in perplexity between our written (2.35) and spoken (2.25) test sets and the presentations given by the AAC user (2.60). The public test sets also produced lower average perplexities (2.28 for ALS bank and 2.22 for COMM2). While we hoped our data selection approach using diverse text from mobile forum messages would allow our adapted models to perform well on both simple everyday communications and more complex planned written communications, at least for this one user, this did not work. As a measure of text complexity, we calculated the number of characters per sentence. The AAC user’s sentences were on average 94.4 characters per sentence compared to the much shorter sentences in the written (31.4) and spoken (39.9) test sets.
The byte LLM provided the fastest performance and was 5.1 times faster than our search algorithm and 1.3 times faster than using a classification layer.